Sorry for the bad english, i dont know chemistry so well in english.

There is 25ml of hydrochloric acid in a volumetric flask, and it's concentration is 0.125mol/l. Then you add 175ml of water, so the liquid becomes 200ml in total. What is the concentration now, when there's 175ml of water added to the flash?
1
Expert's answer
2020-04-03T11:49:39-0400

n = C*V

n = 0.125*0.025 = 0.003125 mol

New volume is 200 mL or 0.2 L

New C = 0.003125 / 0.2 = 0,015625 or 0.016 M.